State the ordered pair that satisfies the equation 2x-5y=-1
a (2,4/5)
b (0,3/5)
c (1,4/5)
d (1,3/5)

Answer by checkley77(12569)

2x-5y=-1
ans. (1,3/5)
Proof:
2*1-5*3/5=-1
2-15/5=-1
2-3=-1
-1=-1